OpenStudy (anonymous):

A newly discovered element, X, has two naturally occurring isotopes. 93.5 percent of the sample is an isotope with a mass of 268.9 u, and 6.5 percent of the sample is an isotope with a mass of 269.9 u. What is the weighted average atomic mass for this element? 268.5 u 269.0 u 269.4 u 269.8 u

OpenStudy (aaronq):

a weighted average takes into account the abundance of each component. \(m_{average}=\Sigma (mass_i*Isotope_{abundance}\) \(m_{average}=(268.9amu)*(0.935)+(269.9amu)*(0.065)\)

OpenStudy (aaronq):

ps. i forgot to close the bracket

OpenStudy (anonymous):

So what would be the answer? Im so lost . @aaronq

OpenStudy (aaronq):

you just need to multiply each pair and then add them together (268.9)*(0.935)+(269.9)*(0.065) heres a calculator http://web2.0calc.com

OpenStudy (asib1214):

atomic mass= % abundance of isotope 1 (mass of isotope1) + % abundance of isotope 2 (mass of isotope 2) +.... Solution= 93.5% (268.9u) + 6.5% (269.9u) = 251.42+17.54 = 268.96u

OpenStudy (aaronq):

round up 268.965 = 268.97

OpenStudy (asib1214):

yep 268.965= 268.97